# Evaluating the Impacts of Swapping on the US Decennial Censusthanks: This work was supported in part by Sloan Foundation grant G-2017-9890, a Google Research Scholar award, and the Center for Research on Computation and Society (CRCS), Harvard

### v1

@@ 4.3 Quantifying the Variance of Swapping
In addition to biasing quantities of interest, because they are randomized algorithms, both swapping and TopDown introduce variance into their outputs. Here, we compare the variance they introduce into block-level counts by race. Our techniques draw inspiration from Kenny et al. (2023b), who establish estimators for the bias and variance introduced by TopDown. Bailie et al. (2023) also compare swapping at different rates against TopDown, but with one key difference: while we compare the data accuracy of swapping and TopDown, they compare the formal privacy guarantees of swapping and ToyDown. Though we do not study formal privacy guarantees, our study of variance is motivated by its connection to privacy. Noise is necessary to ensure privacy Dinur and Nissim (2003), and the variance of a statistical process is a measure of how much noise it introduces.
@@
$=2\operatorname*{\mathbb{E}}\left[\varepsilon_{\gamma}^{r}(\mathbf{X}_{\text{Swapped}})^{2}\right]-2\operatorname*{\mathbb{E}}\left[\varepsilon_{\gamma}^{r}(\mathbf{X}_{\text{Swapped}})\right]^{2}$ ($\varepsilon_{\gamma}^{r}(\mathbf{X}_{\text{Swapped}})$, $\varepsilon_{\gamma}^{r}(\mathbf{X}_{\text{Swapped}}^{\prime})$ i.i.d.)
@@
$\widehat{V}_{\text{TopDown}}$ $\triangleq\frac{1}{2|\mathcal{B}_{s}||\mathcal{R}|}\sum_{\gamma\in\mathcal{B}_{s}}\sum_{r\in\mathcal{R}}\left(c_{\gamma}^{r}(\mathbf{X}_{\text{TopDown2021}})-c_{\gamma}^{r}(\mathbf{X}_{\text{TopDown2023}})\right)^{2}.$ (1)
@@
Figure 4: The variance of swapping at different swap rates. Error bars show the minimum and maximum of 5 runs of our estimator. The horizontal dotted lines show the variance estimate for TopDown. Note that we do not characterize the error in our estimate for TopDown. Plots for two more states are in Figure 9.
Given the connection between variance and privacy described above, these observations suggest that our swapping algorithm with realistic swap rates likely has lower variance, and therefore lower privacy, than TopDown.
@@
$\widehat{V}_{\text{ToyDown}}(\epsilon)\triangleq\frac{1}{2|\mathcal{B}_{s}||\mathcal{R}|}\sum_{\gamma\in\mathcal{B}_{s}}\sum_{r\in\mathcal{R}}(c_{\gamma}^{r}(\mathbf{X}_{\text{ToyDown},\epsilon})-c_{\gamma}^{r}(\mathbf{X}_{\text{ToyDown},\epsilon}^{\prime}))^{2}.$
@@ 4.4 Effect of Swapping on Racial Entropy
We measure the effect of swapping on racial entropy, which we use as a notion of diversity, in Alabama. We define the racial entropy of a geographic area to be the entropy of the empirical distribution of race in that area; i.e., for a geographic area $\gamma$,
@@
Because swapping specifically targets rare households and moves them to a different tract, we might expect it to have a homogenizing effect, resulting in reduced tract-level racial entropy. Surprisingly, we find that, on average, swapping increases the racial entropy of each tract. Table 8 in Appendix C shows the average entropy of tracts before and after swapping; the average entropy increases from unswapped data to swapped data at a 2% rate and again from swapped data at a 2% rate to swapped data at a 10% rate.
@@
    As discussed in Appendix B, household size is a non-racial factor to swapping, and it turns out to be quite influential in determining which households are swapped.
@@
    Households may be swapped not because they have uncommon races, but because their exact composition of races is uncommon. Table 1 shows that mixed race households are over-represented among the households involved in swapping.
@@
Overall, we find that the entropy increasing effects of swapping outweigh the entropy decreasing effects, as shown in Figure 5.
@@
Figure 5: The effect of swapping on racial entropy at the tract level for a 10% swap rate in Alabama. The left panel shows the effect of each step of swapping on entropy. The right panel shows the overall effect. Corresponding plots for 2% and 10% swap rates in other states are in Figures 15, 16, 17, 18, and 19 in Appendix C.
@@ 5.1 Voting
The estimation of voting patterns by race is crucial for applications such as enforcing the Voting Rights Act requirement that voting districts are drawn in a way that allows minority voters to elect their preferred candidates (Freedman et al., 1991). Given many sets of marginal counts of votes and racial groups (one for each voting district or precinct), two common methods are used to infer voting preferences: ecological regression (ER) and ecological inference (EI). Following (Cohen et al., 2021), we focus on ER.
@@
Like (Cohen et al., 2021), we observe that weighting ER by the population of the voting districts reduces the effect of ToyDown to almost nothing. In contrast, weighting does not appear to reduce the effect of swapping in the same way. The result of ER with weighting is shown in Figure 21 in Appendix C. This suggests that interventions that mitigate errors due to differential privacy may not hold for swapping, and perhaps vice versa. Our results indicate that the effects of swapping can differ substantially from those of differential privacy.
@@ 5.2 Social science
With our methods, we can carry out an analogous analysis on the impact of swapping, shown in Figure 7. We replicate the qualitative observations of Mueller and Santos-Lozada (2022) and make new observations about swapping. [^6] Our results are based only on Alabama, Nevada, Texas, Vermont, and Wisconsin, while the results in (Mueller and Santos-Lozada, 2022) are based on the whole U.S. This is because we have not run our swapping algorithm on the whole U.S. for the sake of saving compute time. In principle, it would be possible for us to swap the whole country. We conclude that, while differential privacy indeed introduces problematic relative errors in small racial groups in rural areas, so does swapping. Even at a 2% swap rate (which we believe to be conservative), the relative errors introduced by swapping in counts of small racial groups in rural areas are as great or greater than those introduced by TopDown. [^7] Note that, to match (Mueller and Santos-Lozada, 2022), we actually compare TopDown to the released 2010 data, which is swapped (this corresponds to the arrow marked “Panel D” in Figure 1). As mentioned before, this is a comparison between two different processes applied to the same underlying data, not a direct comparison between the underlying data and the processed data. Note that total population counts are not affected by swapping, so the errors introduced by swapping are always zero (compared to small, but nonzero, errors introduced by TopDown).
@@ Appendix A Additional Swapping Implementation Details
  • We know that the Census Bureau assigns each household to a tier from 1 (least risk) to 4 (greatest risk), where tier 4 households are swapped with probability 1 [Steel and Zayatz, 2003]. To assign these tiers, we sort the households by reidentification risk (using the method above to determine reidentification risk) and split the list into tiers. We set the distribution of tier assignments such that (1) there are twice as many tier 3 households as tier 4 households, and three times as many tier 2 households as tier 4 households and (2) in expectation, we will swap all of the tier 4 households and half of the tier 3 households by the time we make the desired number of swaps (assuming the standard swap probabilities described below). For reference, using a $10\%$ swap rate, this means that $6.25\%$ of the households are in tier 4, $12.5\%$ are in tier 3, $18.75\%$ are in tier 2, and $62.5\%$ are in tier 1.888Since we iterate through households in order of their tier, households in tier 1 or 2 are very unlikely to be targeted to be swapped, but still have a chance of being involved in a swap as the swap partner.
@@ Appendix B Additional Analyses of Swapping
Each swap involves two households: the household that was targeted for swapping (because of its relatively high risk for reidentification) and the household that is selected as the target’s partner. In Table 1, we show the races of the households that are chosen as swap targets with a 10% swap rate in Alabama. In Table 2, we show the distribution of the race of the partner household, for a given type of target household. Corresponding tables for a 2% swap rate in Alabama are Tables 5 and 6 in Appendix C.

### v2

@@ 4.3 Quantifying the Variance of Swapping
In addition to biasing quantities of interest, because they are randomized algorithms, both swapping and TopDown introduce variance into their outputs. Here, we compare the variance they introduce into block-level counts by race. Our techniques draw inspiration from Kenny et al. (2023b), who establish estimators for the bias and variance introduced by TopDown. Bailie et al. (2023) also compare swapping at different rates against TopDown, but with one key difference: while we compare the data accuracy of swapping and TopDown, they compare the formal privacy guarantees of swapping and ToyDown. Though we do not study formal privacy guarantees, our study of variance is motivated by its connection to privacy. Noise is necessary to ensure privacy (Dinur and Nissim, 2003), and the variance of a statistical process is a measure of how much noise it introduces.
@@
$=2\operatorname*{\mathbb{E}}\left[\varepsilon_{\gamma}^{r}(\mathbf{X}_{\text{Swapped}})^{2}\right]-2\operatorname*{\mathbb{E}}\left[\varepsilon_{\gamma}^{r}(\mathbf{X}_{\text{Swapped}})\right]^{2}$ ($\varepsilon_{\gamma}^{r}(\mathbf{X}_{\text{Swapped}})$, $\varepsilon_{\gamma}^{r}(\mathbf{X}_{\text{Swapped}}^{\prime})$ are i.i.d.)
@@
$\widehat{V}_{\text{TopDown}}$
$\triangleq\frac{1}{2|\mathcal{B}_{s}||\mathcal{R}|}\sum_{\gamma\in\mathcal{B}_{s}}\sum_{r\in\mathcal{R}}\left(c_{\gamma}^{r}(\mathbf{X}_{\text{TopDown2021}})-c_{\gamma}^{r}(\mathbf{X}_{\text{TopDown2023}})\right)^{2}.$
@@
Figure 4: The variance of swapping at different swap rates. Error bars show the minimum and maximum of 5 runs of our estimator. The horizontal dotted lines show the variance estimate for TopDown. Note that we do not characterize the error in our estimate for TopDown. Plots for two more states appear in Figure 9.
Given the connection between variance and privacy described above, these observations suggest that our swapping algorithm with realistic swap rates likely has lower variance, and therefore lower privacy, than TopDown. This observation is consistent with statements made by Abowd (2021).
@@
$\widehat{V}_{\text{ToyDown}}(\epsilon)$
$\triangleq\frac{1}{2|\mathcal{B}_{s}||\mathcal{R}|}\sum_{\gamma\in\mathcal{B}_{s}}\sum_{r\in\mathcal{R}}(c_{\gamma}^{r}(\mathbf{X}_{\text{ToyDown},\epsilon})-c_{\gamma}^{r}(\mathbf{X}_{\text{ToyDown},\epsilon}^{\prime}))^{2}.$
@@ 4.4 Effect of Swapping on Racial Entropy
Here, we consider how swapping impacts the racial distribution of the population. We use entropy as our primary measure of interest here. Higher racial entropy would indicate that populations are more geographically segregated. Here, we focus our analysis on Alabama. Formally, define the racial entropy of a geographic area to be the entropy of the empirical distribution of race in that area; i.e., for a geographic area $\gamma$,
@@
Because swapping specifically targets rare households and moves them to a different tract, we might expect it to have a homogenizing effect, resulting in reduced tract-level racial entropy. Surprisingly, we find that, on average, swapping increases the racial entropy of each tract. The average entropy increases from unswapped data to swapped data at a 2% rate and again from swapped data at a 2% rate to swapped data at a 10% rate. Table 9 shows the average entropy of tracts before and after swapping.
State
Average tract entropy
before swapping
Average tract entropy
after 2% swapping
Average tract entropy
after 10% swapping
AL 0.598 0.603 0.624
WI 0.431 0.434 0.447
TX 0.807 0.810 0.819
NV 0.997 0.999 1.006
VT 0.232 0.233 0.241
@@
    Household size is a non-racial factor to swapping, and it turns out to be quite influential in determining which households are swapped.
@@
    Households may be swapped not because they have uncommon races, but because their exact composition of races is uncommon. One reflection of this is that mixed race households are over-represented among the households involved in swapping.
@@
Figure 5: The effect of swapping on racial entropy at the tract level for a 10% swap rate in Alabama. The left panel shows the effect of each step of swapping on entropy. The right panel shows the overall effect. Corresponding plots for 2% and 10% swap rates in other states appear in Figures 15, 16, 17, 18 and 19 in Appendix C.
@@ 5.1 Voting
The estimation of voting patterns by race is crucial for applications such as enforcing the Voting Rights Act requirement that voting districts are drawn in a way that allows minority voters to elect their preferred candidates (Freedman et al., 1991). Given many sets of marginal counts of votes and racial groups (one for each voting district or precinct), two common methods are used to infer voting preferences: ecological regression (ER) and ecological inference (EI). Following Cohen et al. (2021), we focus on ER.
@@
Like Cohen et al. (2021), we observe that weighting ER by the population of the voting districts reduces the effect of ToyDown to almost nothing. In contrast, weighting does not appear to reduce the effect of swapping in the same way. The result of ER with weighting is shown in Figure 21 in Appendix C. This suggests that interventions that mitigate errors due to differential privacy may not hold for swapping, and perhaps vice versa. Our results indicate that the effects of swapping can differ substantially from those of differential privacy.
@@ 5.2 Social science
With our methods, we can carry out an analogous analysis on the impact of swapping, shown in Figure 7. We replicate the qualitative observations of Mueller and Santos-Lozada (2022) and make new observations about swapping. [^7] Our results are based only on Alabama, Nevada, Texas, Vermont, and Wisconsin, while the results in Mueller and Santos-Lozada (2022) are based on the whole U.S. While we limit our analyses to those states for computational reasons, in principle, it would be possible to use our methods to swap the whole country. We conclude that, while differential privacy indeed introduces problematic relative errors in small racial groups in rural areas, so does swapping. Even at a 2% swap rate (which we believe to be conservative), the relative errors introduced by swapping in counts of small racial groups in rural areas are as great or greater than those introduced by TopDown. [^8] Note that, to match Mueller and Santos-Lozada (2022), we actually compare TopDown to the released 2010 data, which is swapped (this corresponds to the arrow marked “Panel D” in Figure 1). As mentioned before, this is a comparison between two different processes applied to the same underlying data, not a direct comparison between the underlying data and the processed data. Note that total population counts are not affected by swapping, so the errors introduced by swapping are always zero (compared to small, but nonzero, errors introduced by TopDown).
@@ Appendix A Additional Swapping Implementation Details
  • We know that the Census Bureau assigns each household to a tier from 1 (least risk) to 4 (greatest risk), where tier 4 households are swapped with probability 1 [Steel and Zayatz, 2003]. To assign these tiers, we sort the households by reidentification risk (using the method above to determine reidentification risk) and split the list into tiers. We set the distribution of tier assignments such that (1) there are twice as many tier 3 households as tier 4 households, and three times as many tier 2 households as tier 4 households and (2) in expectation, we will swap all of the tier 4 households and half of the tier 3 households by the time we make the desired number of swaps (assuming the standard swap probabilities described below). For reference, using a $10\%$ swap rate, this means that $6.25\%$ of the households are in tier 4, $12.5\%$ are in tier 3, $18.75\%$ are in tier 2, and $62.5\%$ are in tier 1.999Since we iterate through households in order of their tier, households in tier 1 or 2 are very unlikely to be targeted to be swapped, but still have a chance of being involved in a swap as the swap partner.
@@ Appendix B Additional Analyses of Swapping
Each swap involves two households: the household that was targeted for swapping (because of its relatively high risk for reidentification) and the household that is selected as the target’s partner. In Table 2, we show the races of the households that are chosen as swap targets with a 10% swap rate in Alabama. In Table 3, we show the distribution of the race of the partner household, for a given type of target household. Corresponding tables for a 2% swap rate in Alabama are Tables 6 and 7 in Appendix C.